A Chemist makes a standard by diluting 10.408 mg Standard to 100 mill in a volumetric flask the standard bottle reads for quantitive applicant use a value of 0.997 MG standard per MG of material on an Aziz basis the chemist makes a sample by dissolving 10 tablets in 1000 ML volumetric flask when the analyst was complete the following absorbance were obtained average of all standard injections equals 5820178 average of duplicate sample injections equals 5600881


1
Expert's answer
2021-01-29T08:31:32-0500

mg/tablet = "\frac{Abs.sample×wgt.sample×purity.std×volume.sample}{Abs.std×volume.std×number.of.tablets}"

Abs sample = 5600881

wgt sample = 10.408mg

Purity std = 0.997mg

Volume sample = 100ml

Abs std = 5820178

Volume std = 1000ml

Number of tablet = 10 tablets

= "\frac{5600881×10.408×0.997×100}{5820178×1000×10}" × 100


= 0.0998579211 × 100

= 9.98579211 %